40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 9471     Entry time: Sat Dec 14 02:51:47 2013
Author: Den 
Type: Update 
Category: LSC 
Subject: locking activity 

I had a look on x,y arms stabilization using ALS. Input green beam was misaligned and I was loosing 00 every few minutes. I vent on the floor and realigned green beams.

YARM alignemt was smooth - transmission increased from 0.4 to 0.85 with PSL shutter off.

XARM was tough. Steering mirrors did not have any derivatives when transmission power was 0.5. I walked the beam with piezos but got only 0.55. It seems that the input beam is mismatched to the cavity. When the transmission was 1 last time? Does anyone have a model of the xend table to compute mode matching?

Input green alignent was improved and I could keep arms stabilized for periods of ~30min - 1 hour. Still not forever.

I noticed that ALS_XARM and ALS_YARM servos have limiters of 6000 and control signal had high frequency components that were not rolled off as shown on the plot "ETMY_DRIVE". I have added a low pass filter that reduced RMS by factor of 5 and took 7 degrees of phase at UGF=150 Hz. Now margin is 33 degrees.

Then I excited ETMY longitudinally at 100 Hz and measured first and second harmonics of the YARM RIN. I got total DC offset of 0.3 nm. This means significant length coupling to RIN. First of all, "scan arm" script does not tune the offset very precise. I guess it looks at DC power, checks when cavity passes through symmetrical points of the resonance and takes the average. It is also useful to look at POX/POY and confirm that average is 0. Plot "ALS_RIN" shows comparison of YARM power fluctuations when it is locked using IR and stabilized using ALS. By manually correcting the offset I could reduce length coupling into RIN, coherence was ~0.1.

Cavity RMS motion also couples length to RIN. Plot "ALS_IR" shows YARM error signal. I also looked at POY signal (LSC-YARM_IN1) as an OOL sensor. At low frequencies POY sees only IMC length fluctuations converted to frequency. I have engaged MCL path and ALS error and LSC error signals overlaped. Cavity RMS motion is measured to be 200 pm.

Attachment 1: ETMY_DRIVE.pdf  46 kB  | Hide | Hide all | Show all
ETMY_DRIVE.pdf
Attachment 2: ALS_RIN.pdf  45 kB  | Show | Hide all | Show all
Attachment 3: ALS_IR.pdf  62 kB  | Show | Hide all | Show all
ELOG V3.1.3-